# On a B-field Transform of Generalized Complex Structures over Complex Tori

### v1

@@ 5.3 A deformation of objects of $Fuk_{\rm sub}(\check{X}^{n})$ and the homological mirror symmetry
As mentioned in the beginning of section 5, we assume $\tau\in M(n;\mathbb{Z})$ in this subsection. Here, under the assumption $\tau\in M(n;\mathbb{Z})$, we discuss the deformation of objects $\mathscr{L}_{(s,a,q)}^{\nabla}\in Fuk_{\rm sub}(\check{X})$ associated to the deformation from $\check{X}^{n}$ to $\check{X}_{\mathcal{G}_{\tau}^{\nabla}}^{n}$, and confirm that such deformed objects and twisted holomorphic line bundles with twisted integrable connections $E_{(s,a,q)}^{\nabla}(\tau)$ are mirror dual to each other. This result can be regarded as a part of the statement of the homological mirror symmetry conjecture for $(X_{\mathcal{G}_{\tau}^{\nabla}}^{n},\check{X}_{\mathcal{G}_{\tau}^{\nabla}}^{n})$, and it is given in Proposition 5.7.
@@
We consider the deformation $\mathscr{L}_{(s,a,q)}^{\nabla}(\tau)$ of a given arbitrary object $\mathscr{L}_{(s,a,q)}^{\nabla}$, namely, we discuss the condition such that $\mathscr{L}_{(s,a,q)}^{\nabla}(\tau)$ gives an object of the Fukaya category over $\check{X}_{\mathcal{G}_{\tau}^{\nabla}}^{n}$ later similarly as in subsection 5.1. It is natural to consider that $L_{(s,a)}$ is not deformed and $\mathcal{L}_{(s,a,q)}^{\nabla}$ is only deformed depending on $\tau$ since $\omega^{\vee}$ is preserved and $B^{\vee}$ is twisted by using $\tau$ in the deformation from $\check{X}^{n}$ to $\check{X}_{\mathcal{G}_{\tau}^{\nabla}}^{n}$, so as candidates of objects of the Fukaya category over $\check{X}_{\mathcal{G}_{\tau}^{\nabla}}^{n}$, we focus on $n$-dimensional submanifolds $L_{(s,a)}$ in $\check{X}_{\mathcal{G}_{\tau}^{\nabla}}^{n}$ similarly as in the case of $Fuk_{\rm sub}(\check{X}^{n})$. Below, we discuss the deformation of $\mathcal{L}_{(s,a,q)}^{\nabla}$. Let us consider a smooth complex line bundle $\mathcal{L}_{(s,a,q)}^{\tau}$ on $L_{(s,a)}$ which is defined by a transition function expressed locally as
@@
$\psi\left(\begin{array}[]{cccccccc}x^{1}\\ \vdots\\ x^{i}+1\\ \vdots\\ x^{n}\\ s^{1}(\check{x})+a_{1i}\\ \vdots\\ s^{n}(\check{x})+a_{ni}\end{array}\right)=\mathrm{exp}\left(\pi\mathbf{i}\tau_{i}^{t}\check{x}\right)\psi\left(\begin{array}[]{cccccccc}x^{1}\\ \vdots\\ x^{i}\\ \vdots\\ x^{n}\\ s^{1}(\check{x})\\ \vdots\\ s^{n}(\check{x})\end{array}\right)$
@@
for each $i=1$, $\cdots$, $n$, where $\psi$ is a locally defined smooth section of $\mathcal{L}_{(s,a,q)}^{\tau}$. We further consider a flat connection on $\mathcal{L}_{(s,a,q)}^{\tau}$ which is expressed locally as
$\nabla_{\mathcal{L}_{(s,a,q)}^{\tau}}:=d-\pi\mathbf{i}\check{x}^{t}\tau d\check{x}-2\pi\mathbf{i}q^{t}d\check{x},$
@@
is a constant vector. It is easy to check that the local expression of the curvature form $\Omega_{\nabla_{\mathcal{L}_{(s,a,q)}^{\tau}}}:=\nabla_{\mathcal{L}_{(s,a,q)}^{\tau}}\circ\nabla_{\mathcal{L}_{(s,a,q)}^{\tau}}$ of $\nabla_{\mathcal{L}_{(s,a,q)}^{\tau}}$ is given by
$\Omega_{\nabla_{\mathcal{L}_{(s,a,q)}^{\tau}}}=-\pi\mathbf{i}d\check{x}^{t}\tau d\check{x},$
and this implies that $\tau\in M(n;\mathbb{Z})$ determines the first Chern character $ch_{1}(\mathcal{L}_{(s,a,q)}^{\tau})$ of $\mathcal{L}_{(s,a,q)}^{\tau}$:
$ch_{1}(\mathcal{L}_{(s,a,q)}^{\tau})=\sum_{1\leq i<j\leq n}\tau_{ij}dx^{i}\wedge dx^{j}.$
Hereafter, for a smooth complex line bundle $\mathcal{L}_{(s,a,q)}^{\tau}$ and a flat connection $\nabla_{\mathcal{L}_{(s,a,q)}^{\tau}}$ on $\mathcal{L}_{(s,a,q)}^{\tau}$, we set
@@
and let us denote a pair of an $n$-dimensional submanifold $L_{(s,a)}$ in $\check{X}_{\mathcal{G}_{\tau}^{\nabla}}^{n}$ and a smooth complex line bundle with a flat connection $\mathcal{L}_{(s,a,q)}^{\nabla}(\tau)$ by
@@
Now, we present the following proposition.
@@
For a given locally defined smooth function $s$ satisfying the relations (18) and a given constant vector $q\in\mathbb{R}^{n}$, the pair $\mathscr{L}_{(s,a,q)}^{\nabla}(\tau)$ gives an object of the Fukaya category over $\check{X}_{\mathcal{G}_{\tau}^{\nabla}}^{n}$ if and only if
holds.
@@
According to [9, Definition 1.1], we may consider the following conditions:
$L_{(s,a)}\ \mathrm{becomes}\ \mathrm{a}\ \mathrm{Lagrangian}\ \mathrm{submanifold}\ \mathrm{in}\ \check{X}_{\mathcal{G}_{\tau}^{\nabla}}^{n}.$ (27)
$\mathcal{L}_{(s,a,q)}^{\tau}\to L_{(s,a)}\ \mathrm{with}\ \nabla_{\mathcal{L}_{(s,a,q)}^{\tau}}\ \mathrm{satisfies}\ \Omega_{\nabla_{\mathcal{L}_{(s,a,q)}^{\tau}}}=\Bigl{.}-\mathbf{i}B_{\tau}^{\vee}\Bigr{|}_{L_{(s,a)}}.$ (28)

### v2

@@ 5.4 A deformation of objects of $Fuk_{\rm sub}(\check{X}^{n})$ and the homological mirror symmetry
The purpose of this subsection is to explain the details of the proposal which is mentioned in subsection 5.3, namely, to discuss the deformation of objects $\mathscr{L}_{(s,a,q)}^{\nabla}\in Fuk_{\rm sub}(\check{X}^{n})$ which is compatible with the counterpart in the mirror dual complex geometry side.
Let us first define a flat gerbe $\check{\mathcal{G}}_{\tau_{\rm frac}}^{\nabla}$ over the complexified symplectic torus
$\check{X}^{n}(\tau_{\mathbb{Z}}):=\Bigl{(}\mathbb{R}^{2n}/\mathbb{Z}^{2n},\ \omega^{\vee}-\mathbf{i}(B^{\vee}+\pi d\check{x}^{t}\tau_{\mathbb{Z}}d\check{x})\Bigr{)}$
depending on $\tau_{\rm frac}\in M(n;\mathbb{R})$ as follows. We fix an $\epsilon>0$ small enough and take an open covering $\{\check{O}_{m}^{l}:=\check{O}_{m_{1}\cdots m_{n}}^{l_{1}\cdots l_{n}}\}_{(l;m)\in I}$ of $\check{X}^{n}$ which is obtained by replacing $\left(\begin{array}[]{cc}x\\ y\end{array}\right)$ with $\left(\begin{array}[]{cc}\check{x}\\ \check{y}\end{array}\right)$, i.e.,
$\check{O}_{m_{1}\cdots m_{n}}^{l_{1}\cdots l_{n}}:=\biggl{\{}\left(\begin{array}[]{ccc}\check{x}\\ \check{y}\end{array}\right)\in\check{X}^{n}\ \Bigl{.}\Bigr{|}$ $\frac{l_{j}-1}{3}-\epsilon<x^{j}<\frac{l_{j}}{3}+\epsilon,$
and in particular, we can also regard this $\{\check{O}_{m}^{l}\}_{(l;m)\in I}$ as an open covering of $\check{X}^{n}(\tau_{\mathbb{Z}})$. Let us consider the trivial gerbe $\check{\mathcal{G}}(I,\mathcal{O},\xi)$ on $\check{X}^{n}(\tau_{\mathbb{Z}})$ (similarly as in the case of $\mathcal{G}(I,\mathcal{O},\xi)$, we use the notations $\mathcal{O}$ and $\xi$ here):
Similarly, a 0-connection ˇ$\check{\nabla}_{\tau_{\rm frac}}:=\{^{\tau_{\rm frac}}\!\check{\nabla}_{mm^{\prime}}^{ll^{\prime}}:=d-\mathbf{i}^{\tau_{\rm frac}}\!\check{\omega}_{mm^{\prime}}^{ll^{\prime}}\}_{(l;m),(l^{\prime};m^{\prime})\in I}$ over $\check{\mathcal{G}}(I,\mathcal{O},\xi)$ is also defined by replacing $\tau$ and $x$ with $\tau_{\rm frac}$ and $\check{x}$, respectively, namely, a 1-form ${}^{\tau_{\rm frac}}\!\check{\omega}_{mm^{\prime}}^{ll^{\prime}}$ on the open set $\check{O}_{mm^{\prime}}^{ll^{\prime}}$ is defined by
${}^{\tau_{\rm frac}}\!\check{\omega}_{mm^{\prime}}^{ll^{\prime}}=\begin{cases}\pi(\tau_{\rm frac})_{j}^{t}d\check{x}&(l_{j}=1,l_{j}^{\prime}=3,l_{h}=l_{h}^{\prime}(h\not=j),m=m^{\prime})\\ \pi(\tau_{\rm frac})_{j}^{t}d\check{x}&(l_{j}=1,l_{j}^{\prime}=3,l_{h}=l_{h}^{\prime}(h\not=j),m_{k}=1,m_{k}^{\prime}=3,m_{i}=m_{i}^{\prime}(i\not=k))\\ \pi(\tau_{\rm frac})_{j}^{t}d\check{x}&(l_{j}=1,l_{j}^{\prime}=3,l_{h}=l_{h}^{\prime}(h\not=j),m_{k}=3,m_{k}^{\prime}=1,m_{i}=m_{i}^{\prime}(i\not=k))\\ 0&(\mathrm{otherwise}),\end{cases}$
where
$(\tau_{\rm frac})_{j}:=\left(\begin{array}[]{ccc}(\tau_{\rm frac})_{j1}\\ \vdots\\ (\tau_{\rm frac})_{jn}\end{array}\right)\in\mathbb{R}^{n},$
$j$, $k=1$, $\cdots$, $n$, and we assume the condition ${}^{\tau_{\rm frac}}\!\check{\omega}_{m^{\prime}m}^{l^{\prime}l}=-^{\tau_{\rm frac}}\!\check{\omega}_{mm^{\prime}}^{ll^{\prime}}$. Moreover, although we drop the index $(l;m)\in I$ for simplicity, the family
$\Bigl{\{}-\pi\mathbf{i}d\check{x}^{t}\tau_{\rm frac}d\check{x}\Bigr{\}}_{(l;m)\in I}$
gives a 1-connection over $\check{\mathcal{G}}(I,\mathcal{O},\xi)$ which is compatible with the 0-connection ˇ$\check{\nabla}_{\tau_{\rm frac}}$. We denote the trivial gerbe $\check{\mathcal{G}}(I,\mathcal{O},\xi)$ with the flat connection (ˇ$(\check{\nabla}_{\tau_{\rm frac}},\{-\pi\mathbf{i}d\check{x}^{t}\tau_{\rm frac}d\check{x}\}_{(l;m)\in I})$ by $\check{\mathcal{G}}_{\tau_{\rm frac}}^{\nabla}$:
$\check{\mathcal{G}}_{\tau_{\rm frac}}^{\nabla}:=\Bigl{(}\check{\mathcal{G}}(I,\mathcal{O},\xi),\ \check{\nabla}_{\tau_{\rm frac}},\ \Bigl{\{}-\pi\mathbf{i}d\check{x}^{t}\tau_{\rm frac}d\check{x}\Bigr{\}}_{(l;m)\in I}\Bigr{)}.$
By using this flat gerbe $\check{\mathcal{G}}_{\tau_{\rm frac}}^{\nabla}$, similarly as in the deformation $X_{\mathcal{G}_{\tau}^{\nabla}}^{n}$ of $X^{n}$ by $\mathcal{G}_{\tau}^{\nabla}$, we can consider the deformation $\check{X}^{n}(\tau_{\mathbb{Z}})_{\check{\mathcal{G}}_{\tau_{\rm frac}}^{\nabla}}$ of $\check{X}^{n}(\tau_{\mathbb{Z}})$ by $\check{\mathcal{G}}_{\tau_{\rm frac}}^{\nabla}$, and $\check{X}_{\mathcal{G}_{\tau}^{\nabla}}^{n}$ defined in section 4 can be identified with this $\check{X}^{n}(\tau_{\mathbb{Z}})_{\check{\mathcal{G}}_{\tau_{\rm frac}}^{\nabla}}$. By concerning this fact, hereafter, we treat $\check{X}^{n}(\tau_{\mathbb{Z}})_{\check{\mathcal{G}}_{\tau_{\rm frac}}^{\nabla}}$ as a mirror partner of $X_{\mathcal{G}_{\tau}^{\nabla}}^{n}$:
$\check{X}_{\mathcal{G}_{\tau}^{\nabla}}^{n}=\check{X}^{n}(\tau_{\mathbb{Z}})_{\check{\mathcal{G}}_{\tau_{\rm frac}}^{\nabla}}.$
@@
Let us discuss the deformation of a given arbitrary object $\mathscr{L}_{(s,a,q)}^{\nabla}\in Fuk_{\rm sub}(\check{X}^{n})$ associated to the deformation from $\check{X}^{n}$ to $\check{X}_{\mathcal{G}_{\tau}^{\nabla}}^{n}$. It is natural to consider that $L_{(s,a)}$ is not deformed and $\mathcal{L}_{(s,a,q)}^{\nabla}$ is only deformed depending on $\tau$ since $\omega^{\vee}$ is preserved and $B^{\vee}$ is twisted by using $\tau$ in the deformation from $\check{X}^{n}$ to $\check{X}_{\mathcal{G}_{\tau}^{\nabla}}^{n}$. Therefore, as candidates of Lagrangian submanifolds, we focus on $n$-dimensional submanifolds $L_{(s,a)}$ in $\check{X}_{\mathcal{G}_{\tau}^{\nabla}}^{n}$ similarly as in the case of $Fuk_{\rm sub}(\check{X}^{n})$. Below, we construct the deformation of $\mathcal{L}_{(s,a,q)}^{\nabla}$. Let us consider an $id$-twisted smooth complex line bundle $\mathcal{L}_{(s,a,q)}^{\tau}\to L_{(s,a)}$ which is defined by a transition function expressed locally as
@@
$\psi\left(\begin{array}[]{cccccccc}x^{1}\\ \vdots\\ x^{i}+1\\ \vdots\\ x^{n}\\ s^{1}(\check{x})+a_{1i}\\ \vdots\\ s^{n}(\check{x})+a_{ni}\end{array}\right)=\mathrm{exp}\left(\pi\mathbf{i}(\tau_{\mathbb{Z}})_{i}^{t}\check{x}\right)\cdot\psi\left(\begin{array}[]{cccccccc}x^{1}\\ \vdots\\ x^{i}\\ \vdots\\ x^{n}\\ s^{1}(\check{x})\\ \vdots\\ s^{n}(\check{x})\end{array}\right)$
@@
for each $i=1$, $\cdots$, $n$, where $\psi$ is a locally defined smooth section of $\mathcal{L}_{(s,a,q)}^{\tau}$, and
$(\tau_{\mathbb{Z}})_{i}:=\left(\begin{array}[]{ccc}(\tau_{\mathbb{Z}})_{i1}\\ \vdots\\ (\tau_{\mathbb{Z}})_{in}\end{array}\right)\in\mathbb{Z}^{n}.$
We further consider a twisted connection on $\mathcal{L}_{(s,a,q)}^{\tau}$ which is expressed locally as
$\nabla_{\mathcal{L}_{(s,a,q)}^{\tau}}:=d-\pi\mathbf{i}\check{x}^{t}\tau_{\mathbb{Z}}d\check{x}-2\pi\mathbf{i}q^{t}d\check{x}-\pi\mathbf{i}\check{x}^{t}\tau_{\rm frac}d\check{x},$
@@
is a constant vector, and $-\pi\mathbf{i}\check{x}^{t}\tau_{\rm frac}d\check{x}$ comes from the deformation of $\check{X}^{n}(\tau_{\mathbb{Z}})$ by $\mathcal{G}_{\tau_{\rm frac}}^{\nabla}$. It is easy to check that the local expression of the curvature form $\Omega_{\nabla_{\mathcal{L}_{(s,a,q)}^{\tau}}}:=\nabla_{\mathcal{L}_{(s,a,q)}^{\tau}}\circ\nabla_{\mathcal{L}_{(s,a,q)}^{\tau}}$ of $\nabla_{\mathcal{L}_{(s,a,q)}^{\tau}}$ is given by
$\Omega_{\nabla_{\mathcal{L}_{(s,a,q)}^{\tau}}}=-\pi\mathbf{i}d\check{x}^{t}\tau_{\mathbb{Z}}d\check{x}.$
At least, this implies that $\tau_{\mathbb{Z}}\in M(n;\mathbb{Z})$ determines the first Chern character $ch_{1}(\mathcal{L}_{(s,a,q)}^{\tau})$ of $\mathcal{L}_{(s,a,q)}^{\tau}$ in the case $\tau\in M(n;\mathbb{Z})$, i.e., $\tau_{\rm frac}=O$:
$ch_{1}(\mathcal{L}_{(s,a,q)}^{\tau})=\frac{1}{2}d\check{x}^{t}\tau_{\mathbb{Z}}d\check{x}.$
Hereafter, for an $id$-twisted smooth complex line bundle $\mathcal{L}_{(s,a,q)}^{\tau}$ and a twisted connection $\nabla_{\mathcal{L}_{(s,a,q)}^{\tau}}$ on $\mathcal{L}_{(s,a,q)}^{\tau}$, we set
@@
and let us denote a pair of an $n$-dimensional submanifold $L_{(s,a)}$ in $\check{X}_{\mathcal{G}_{\tau}^{\nabla}}^{n}$ and an $id$-twisted smooth complex line bundle with a twisted connection $\mathcal{L}_{(s,a,q)}^{\nabla}(\tau)$ by $\mathscr{L}_{(s,a,q)}^{\nabla}(\tau)$:
@@
Then, we obtain the following.
@@
$L_{(s,a)}\ \mathrm{becomes}\ \mathrm{a}\ \mathrm{Lagrangian}\ \mathrm{submanifold}\ \mathrm{in}\ \check{X}_{\mathcal{G}_{\tau}^{\nabla}}^{n}$ (28)
and the condition
if and only if
$\frac{\partial s}{\partial\check{x}}(\check{x})T=\left(\frac{\partial s}{\partial\check{x}}(\check{x})T\right)^{t}.$ (30)